# Cross-lingual Offensive Language Detection: A Systematic Review of Datasets, Transfer Approaches and Challenges

### v1

@@ 7.3.2. Cross-lingual Transfer Performance
CLTL has emerged as a promising solution and has indeed shown effectiveness when addressing data scarcity (Bigoulaeva et al., 2022; Sohn and Lee, 2019). However, offensive language is deeply rooted in the specificity and diversity of language and culture, and its complexity poses significant obstacles to cross-lingual transfer, especially in zero-shot settings (Nozza, 2021; Montariol et al., 2022). Zero-shot transfer approaches to multilingual training often suffer from performance deficiencies when compared to models trained on actual target language data (Pelicon et al., 2021b).
@@ 7.3.4. Poor Model Interpretability
While advanced deep learning-based models have shown good performance in detecting hate speech, they often operate as “black boxes” with a lack of transparency in the decision-making process (Aluru et al., 2021). This opacity makes it hard for humans to understand the underlying reasons for the model performance in a particular language or scenario, and analyse the model errors (Deshpande et al., 2022; Vitiugin et al., 2021). To bridge this interpretability gap, some tools are utilised to explain what models are doing, such as Local Interpretable Model-agnostic Explanations (LIME) (Ribeiro et al., 2016) and SHapley Additive exPlanations (SHAP) (Lundberg and Lee, 2017), and theoretical approaches from cognitive linguistics are also applied, e.g. frame semantics (Vitiugin et al., 2021). A human-in-the-loop paradigm can also help by integrating human expertise into the model’s learning process, but it can be difficult and uncertain for individuals to provide effective feedback to enhance the model performance (Vitiugin et al., 2021).
@@ 7.4. Future Directions
This section outlines promising research directions in this critical research area.
@@ 7.4.1. Dataset Creation
There is a pressing need to collect and annotate comprehensive and balanced datasets, especially in low-resource languages (Husain and Uzuner, 2022; Arango Monnar et al., 2022; Gaikwad et al., 2021; Burtenshaw and Kestemont, 2021). While zero-shot cross-lingual transfer has its limitations, even a small amount of target data can substantially enhance model fine-tuning (R”̈ottger et al., 2022a). To avoid dataset bias, such datasets can span multiple languages, dialects, cultural contexts, and diverse types of offensive language (Ousidhoum et al., 2019). More diverse strategies for data collection, such as sampling from various platforms and user demographics, can yield richer and more annotation-worthy data (R”̈ottger et al., 2022a). Additionally, the creation of domain-specific evaluation datasets, such as XHATE-999 (Glavaš et al., 2020) and Multilingual HateCheck (MHC) (R”̈ottger et al., 2022b), provides a multifaceted and precise perspective for assessing cross-lingual detection.
@@ 7.4.2. Data Annotation
Refining data annotation strategies is instrumental. Instead of large annotating tasks, annotation can be done incrementally, ensuring efficiency and encouraging diverse data collection (R”̈ottger et al., 2022a). In addition, semi-supervised learning methods can efficiently use data and annotator resources and mitigate data scarcity issues. For example, deliberate selection strategies, such as active learning, can enhance annotation performance with fewer target-language entries, mainly annotating data points that the model finds most ambiguous (R”̈ottger et al., 2022a). To further address ambiguity and sensitivity in offensive language, incorporating feedback from human experts can be invaluable (Vitiugin et al., 2021). It would also be better for annotators to follow ethical guidelines and receive appropriate training and support, avoiding potential harm to them.
@@ 7.4.3. Integration of Additional Features
Integrating additional features can also be an effective strategy. Typographic features, including capitals, punctuation, and emojis, offer a language-agnostic and domain-agnostic perspective to offensive language detection. Emojis, associated with emotional expressions, have been highlighted for their ability to help identify online hate due to their shared meanings across languages (Kumar and Sachdeva, 2022; Markov et al., 2021; Corazza et al., 2020). Stylometric features are also robust indicators of offensive language, as the writing style of toxic content can be correlated to social media users (Markov et al., 2021). In addition, domain-specific features are a critical aspect in additional features to assist the knowledge transfer process, especially when addressing linguistic nuances like metaphors, metonymy, and informal expressions. The multilingual lexicon, HurtLex, has been highlighted for its effectiveness in multiple studies (Pamungkas et al., 2020, 2021a; Pamungkas and Patti, 2019; Jiang and Zubiaga, 2021). Since general-purpose multilingual embeddings may not effectively capture some hate-specific patterns, some researchers also utilise hateful word pairs to construct domain-aware or semantic-based representations across languages (Jiang and Zubiaga, 2023; Hahn et al., 2021; Arango et al., 2021). Furthermore, infusing cultural features underscores the significance of language-specific knowledge for geographically sensitive tasks. The creation of culturally-aware models, informed by multidisciplinary studies involving anthropologists and sociologists, can provide a richer understanding of cultural dimensions in hateful content (Zhou et al., 2023). Such work can produce different forms of cultural features, such as switching pattern matrix (Bansal et al., 2020), bilingual language pairs (Jiang and Zubiaga, 2023), cross-cultural similarities (Arango et al., 2021), and social dynamics among users (Nozza, 2021).
@@ 7.4.4. Multilingual Pre-Trained Language Models
Multilingual PLMs can be applied to low-resource target languages without further training, thus bridging the data availability gap between high- and low-resource languages. However, there are significant computational costs during training and fine-tuning due to their large number of parameters (Pikuliak et al., 2021). Future research is geared towards optimising these multilingual PLMs for enhanced efficiency and interpretability. Additionally, two strategies are emerging to improve the generalisability and scalability of multilingual PLMs so that they can scale to handle large datasets or multiple languages simultaneously (Arango et al., 2022). The first emphasises pre-training models using data from relevant sources (such as social platforms and hateful domains), as demonstrated by XLM-T (Montariol et al., 2022) and AbuseXLMR (Gupta et al., 2022). The second strategy focuses on models based on specific low-resource languages or those from similar language families, as seen with MuRIL for Indic languages (Das et al., 2022) and AraBERT for dialects (Husain and Uzuner, 2022).
@@ 7.4.5. Cross-lingual Training Strategies
Researchers are also delving into a number of innovative cross-lingual training strategies. Adaptive training techniques can dynamically adjust to the specificities of different languages and dialects, as well as bridge the gap between offensive language detection and PLMs. By leveraging meta-learning or few-shot learning, models can be trained to quickly adapt to new low-resource languages with few labelled data (Mozafari et al., 2022). Furthermore, collaborative multi-task learning can make models more generalisable in cultural variations and resistant to overfitting by training them and sharing representations on multiple auxiliary tasks simultaneously across different languages or similar NLP tasks (Montariol et al., 2022). The use of adversarial training, where models are trained against adversarial examples, can also enhance their robustness and generalisation across languages. This is especially important for offensive language, as attackers often use subtle language tricks to obfuscate and bypass detection (Shi et al., 2022; Jin et al., 2020; Ryzhova et al., 2022). Moreover, cascade learning can greatly bolster model performance and enhance out-of-domain generalisability by initially training models on source data or more diverse datasets and subsequently fine-tuning the results by simultaneously or progressively incorporating target language data (R”̈ottger et al., 2022a; Pelicon et al., 2021b). Besides, Vitiugin et al. (2021) highlight that human feedback can be of great value to help detect subtleties of abuse and phrases during model training, Ranasinghe and Zampieri (2021c) explore language-specific preprocessing like segmentation in morphologically rich languages such as Arabic and Turkish, and multi-source training, with augmented datasets, is also able to capture the diversities and commonalities of linguistic patterns across languages (Hande et al., 2021a; Markov et al., 2021; Chiril et al., 2019).
@@ 7.4.6. Application of Large Language Models
The emergence of Large Language Models (LLMs), such as GPT-3 (Brown et al., 2020), GPT-4 (OpenAI, 2023), Fine-tuned LAnguage Net (FLAN) (Wei et al., 2021) and Large Language Model Meta AI (LlaMA) (Touvron et al., 2023), has opened up new horizons for cross-lingual offensive language detection, thanks to their remarkable performance (Zhao et al., 2023). Instruction fine-tuned LLMs with prompting have shown promise in detecting offensive language across both mono and multilingual contents without the need for language-specific fine-tuning (Plaza-del arco et al., 2023). The efficacy of LLMs can be further enhanced through well-designed, task-specific prompts (Chen et al., 2022; Han and Tang, 2022; Roy et al., 2023). By incorporating informative task descriptions and input-label demonstrations, these prompts can guide LLMs to achieve superior detection. Furthermore, the inherent multi-tasking capabilities of LLMs allow them to excel not only in offensive language detection but also in some related NLP tasks (Zhang et al., 2023). To address the challenge of data scarcity, LLMs offer a potential solution by synthetically generating hateful texts, especially for languages with limited resources (Hartvigsen et al., 2022). Beyond detection, LLMs can also produce explanations for predictions, enhancing both model performance and interpretability (Wang et al., 2023). While the potential of LLMs in this domain is evident, current research is still in its infancy.
@@ 8. Conclusion
While providing a clear focus and an in-depth discussion on cross-lingual resources and methods, our survey does not cover offensive language datasets, resources and methods outside this scope.
@@ Acknowledgements.
- Alakrot et al. (2018) Azalden Alakrot, Liam Murray, and Nikola S Nikolov. 2018. Dataset construction for the detection of anti-social behaviour in online communication in Arabic. Procedia Computer Science 142 (2018), 174–181.
@@
- Aluru et al. (2021) Sai Saketh Aluru, Binny Mathew, Punyajoy Saha, and Animesh Mukherjee. 2021. A Deep Dive into Multilingual Hate Speech Classification. In Proc. of ECML PKDD, Yuxiao Dong, Georgiana Ifrim, Dunja Mladenić, Craig Saunders, and Sofie Van Hoecke (Eds.). Springer, Cham, 423–439.
@@
- Bosco et al. (2018) Cristina Bosco, Dell’Orletta Felice, Fabio Poletto, Manuela Sanguinetti, Tesconi Maurizio, et al. 2018. Overview of the evalita 2018 hate speech detection task. In Proc. of CEUR Workshop, Vol. 2263. CEUR, 1–9.
@@
- Chen et al. (2022) Yulong Chen, Yang Liu, Li Dong, Shuohang Wang, Chenguang Zhu, Michael Zeng, and Yue Zhang. 2022. AdaPrompt: Adaptive Model Training for Prompt-based NLP. In Findings of EMNLP. ACL, Abu Dhabi, United Arab Emirates, 6057–6068.
@@
- Deng et al. (2022) Jiawen Deng, Jingyan Zhou, Hao Sun, Chujie Zheng, Fei Mi, Helen Meng, and Minlie Huang. 2022. COLD: A Benchmark for Chinese Offensive Language Detection. In Proc. of EMNLP. ACL, Abu Dhabi, United Arab Emirates, 11580–11599.
@@
- Firmino et al. (2021) Anderson Almeida Firmino, Cláudio Souza de Baptista, and Anselmo Cardoso de Paiva. 2021. Using Cross Lingual Learning for Detecting Hate Speech in Portuguese. In Database and Expert Systems Applications, Christine Strauss, Gabriele Kotsis, A. Min Tjoa, and Ismail Khalil (Eds.). Springer, Cham, 170–175.
@@
- Jeong et al. (2022) Younghoon Jeong, Juhyun Oh, Jongwon Lee, Jaimeen Ahn, Jihyung Moon, Sungjoon Park, and Alice Oh. 2022. KOLD: Korean Offensive Language Dataset. In Proc. of EMNLP. ACL, Abu Dhabi, United Arab Emirates, 10818–10833.
@@
- Lundberg and Lee (2017) Scott M Lundberg and Su-In Lee. 2017. A Unified Approach to Interpreting Model Predictions. In Proc. of NeurIPS 30, I. Guyon, U. V. Luxburg, S. Bengio, H. Wallach, R. Fergus, S. Vishwanathan, and R. Garnett (Eds.). Curran Associates, Inc., 4765–4774.
@@
- Markov et al. (2021) Ilia Markov, Nikola Ljubešić, Darja Fišer, and Walter Daelemans. 2021. Exploring Stylometric and Emotion-Based Features for Multilingual Cross-Domain Hate Speech Detection. In Proc. of the WASSA Workshop. ACL, Online, 149–159.
@@
- Mubarak et al. (2021) Hamdy Mubarak, Ammar Rashed, Kareem Darwish, Younes Samih, and Ahmed Abdelali. 2021. Arabic Offensive Language on Twitter: Analysis and Experiments. In Proc. of the WANLP Workshop. ACL, Kyiv, Ukraine (Virtual), 126–135.
@@
- Ranasinghe and Zampieri (2021c) Tharindu Ranasinghe and Marcos Zampieri. 2021c. Multilingual Offensive Language Identification for Low-Resource Languages. ACM Transactions on Asian and Low-Resource Language Information Processing 21, 1, Article 4 (Nov. 2021), 13 pages.
@@
- Rivaling (2021) Bilingual Word Embeddings Rivaling. 2021. Llod-driven bilingual word embeddings rivaling cross-lingual transformers in quality of life concept detection from french online health communities. In Proc. of of SEMANTiCS, Vol. 53. IOS Press, 89.
@@
- Rodríguez et al. (2021) Sebastián E. Rodríguez, Héctor Allende-Cid, and Héctor Allende. 2021. Detecting Hate Speech in Cross-Lingual and Multi-lingual Settings Using Language Agnostic Representations. In Proc. of CIARP, João Manuel R. S. Tavares, João Paulo Papa, and Manuel González Hidalgo (Eds.). Springer, Cham, 77–87.
- Rodríguez-Sánchez et al. (2021) Francisco Rodríguez-Sánchez, Jorge Carrillo-de Albornoz, Laura Plaza, Julio Gonzalo, Paolo Rosso, Miriam Comet, and Trinidad Donoso. 2021. Overview of exist 2021: sexism identification in social networks. Procesamiento del Lenguaje Natural 67 (2021), 195–207.
- Rodríguez-Sánchez et al. (2022) Francisco Rodríguez-Sánchez, Jorge Carrillo-de Albornoz, Laura Plaza, Adrián Mendieta-Aragón, Guillermo Marco-Remón, Maryna Makeienko, María Plaza, Julio Gonzalo, Damiano Spina, and Paolo Rosso. 2022. Overview of exist 2022: sexism identification in social networks. Procesamiento del Lenguaje Natural 69 (2022), 229–240.
@@
- R”̈ottger et al. (2022a) Paul R”̈ottger, Debora Nozza, Federico Bianchi, and Dirk Hovy. 2022a. Data-Efficient Strategies for Expanding Hate Speech Detection into Under-Resourced Languages. In Proc. of EMNLP. ACL, Abu Dhabi, United Arab Emirates, 5674–5691.
- R”̈ottger et al. (2022b) Paul R”̈ottger, Haitham Seelawi, Debora Nozza, Zeerak Talat, and Bertie Vidgen. 2022b. Multilingual HateCheck: Functional Tests for Multilingual Hate Speech Detection Models. In Proc. of the WOAH Workshop. ACL, Seattle, Washington (Hybrid), 154–169.
@@
- Ryzhova et al. (2022) Anastasia Ryzhova, Dmitry Devyatkin, Sergey Volkov, and Vladimir Budzko. 2022. Training Multilingual and Adversarial Attack-Robust Models for Hate Detection on Social Media. Procedia Computer Science 213 (2022), 196–202.
@@
- Stamou et al. (2022) Vivian Stamou, Iakovi Alexiou, Antigone Klimi, Eleftheria Molou, Alexandra Saivanidou, and Stella Markantonatou. 2022. Cleansing & expanding the HURTLEX(el) with a multidimensional categorization of offensive words. In Proc. of the WOAH Workshop. ACL, Seattle, Washington (Hybrid), 102–108.
@@
- Strus̈ et al. (2019) Julia Maria Strus̈, Melanie Siegel, Josef Ruppenhofer, Michael Wiegand, and Manfred Klenner. 2019. Overview of GermEval Task 2, 2019 shared task on the identification of offensive language. In Proc. of KONVENS. München [u.a.], 352 – 363.
@@
- Touvron et al. (2023) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. 2023. Llama: Open and efficient foundation language models. arXiv preprint arXiv:2302.13971 (2023).
@@
- Zampieri et al. (2019) Marcos Zampieri, Shervin Malmasi, Preslav Nakov, Sara Rosenthal, Noura Farra, and Ritesh Kumar. 2019. SemEval-2019 Task 6: Identifying and Categorizing Offensive Language in Social Media (OffensEval). In Proc. of the SemEval Workshop, Jonathan May, Ekaterina Shutova, Aurelie Herbelot, Xiaodan Zhu, Marianna Apidianaki, and Saif M. Mohammad (Eds.). ACL, Minneapolis, Minnesota, USA, 75–86.

### v2

@@ 7.3.2. Cross-lingual Transfer Performance
CLTL has emerged as a promising solution and has shown effectiveness addressing data scarcity (Bigoulaeva et al., 2022; Sohn and Lee, 2019). However, offensive language is deeply rooted in the specificity and diversity of language and culture, and its complexity poses significant obstacles to cross-lingual transfer, especially in zero-shot settings (Nozza, 2021; Montariol et al., 2022). Zero-shot transfer to multilingual training often suffers from performance deficiencies when compared to models trained on actual target language data (Pelicon et al., 2021b). It relies heavily on surface-level lexical and structural features, making them less effective when confronted with cultural contexts or implicit forms of abuse. This limitation has been consistently observed in empirical studies: zero-shot transfer yields significantly weaker results than target-language fine-tuning, even when only small amounts of labelled target data are available (Pelicon et al., 2021b). Moreover, cultural and linguistic variations across communities can further erode transferability, leading to biased or unreliable detection outcomes (Nozza, 2021).
@@ 7.3.4. Poor Model Interpretability
While advanced deep learning-based models have shown good performance in detecting HS, they often operate as “black boxes” with a lack of transparency in the decision-making process (Aluru et al., 2021). This opacity makes it hard for humans to understand the underlying reasons for the model performance in a particular language or scenario, and analyse the model errors (Deshpande et al., 2022; Vitiugin et al., 2021). To bridge this interpretability gap, some tools are utilised to explain what models are doing, such as Local Interpretable Model-agnostic Explanations (LIME) (Ribeiro et al., 2016) and SHapley Additive exPlanations (SHAP) (Lundberg and Lee, 2017), and theoretical approaches from cognitive linguistics are also applied, e.g. frame semantics (Vitiugin et al., 2021). A human-in-the-loop paradigm can also help by integrating human expertise into the model’s learning process, but it can be difficult and uncertain for individuals to provide effective feedback to enhance the model performance (Vitiugin et al., 2021).
@@ 7.4. Future Directions
This section outlines promising research directions and synthesises lessons from prior work.
@@ 7.4.1. Dataset Creation
A recurring finding is that collecting even small amounts of target-language data works well: fine-tuning with just a few hundred annotated examples consistently outperforms pure zero-shot transfer (Röttger et al., 2022a). Conversely, relying solely on zero-shot transfer does not work well, as it leads to significant drops in performance for typologically distant or culturally distinct languages (Pelicon et al., 2021b). Datasets should therefore be designed to cover multiple languages, dialects, and cultural contexts (Ousidhoum et al., 2019), while avoiding narrow topical bias from single platforms (Röttger et al., 2022a). Complementary benchmark resources such as XHATE-999 (Glavaš et al., 2020) and Multilingual HateCheck (MHC) (Röttger et al., 2022b) provide more reliable cross-lingual evaluation.
@@ 7.4.2. Data Annotation
Current evidence suggests that incremental and selective annotation works well: starting with small subsets and iteratively expanding the dataset increases efficiency while avoiding redundant labelling (Röttger et al., 2022a). In contrast, large-scale undirected annotation does not work well, as benefits quickly diminish relative to cost. To advance annotation quality, researchers should recruit diverse annotators and provide culturally contextualised guidelines, which helps mitigate annotation bias. Moreover, active learning and semi-supervised strategies have been shown to optimise annotator effort by focusing on ambiguous or uncertain cases (Röttger et al., 2022a). Expert feedback is particularly valuable for subtle and implicit abuse (Vitiugin et al., 2021), but annotator well-being and ethical safeguards must also be prioritised, providing appropriate training and support to avoid potential harm to them.
@@ 7.4.3. Integration of Additional Features
Studies consistently show that integration of additional features can also be effective. Adding language-agnostic signals such as typographic features (capitalisation, punctuation, emojis) works well (Kumar and Sachdeva, 2022; Markov et al., 2021; Corazza et al., 2020). Emojis, in particular, often carry stable meanings across languages and provide additional emotional cues that complement textual signals. Stylometric features such as function word usage, sentence length, or character-level distributions can also strengthen detection, since they capture writing style cues of toxic content that can be correlated with specific social media users (Markov et al., 2021). Similarly, domain-specific resources such as HurtLex and tailored embeddings work well in capturing hate-related semantics across languages (Pamungkas et al., 2020; Jiang and Zubiaga, 2023). For example, multilingual lexicons of offensive terms have been shown to boost recall by reducing false negatives, while domain-aware embeddings such as sexism-specific representations (SexWEs) provide finer-grained distinctions for subtle or implicit abuse. Relying only on general-purpose multilingual embeddings does not perform well, which often misses culture-specific or implicit abuse patterns (Arango et al., 2021). They may capture general semantic similarity but struggle with figurative expressions, reclaimed slurs, or locally grounded derogatory terms. Over-reliance on generic embeddings risks reinforcing biases by ignoring context-specific linguistic cues. Furthermore, infusing cultural features underscores the significance of language-specific knowledge for geographically sensitive tasks. The creation of culturally-aware models, informed by multidisciplinary studies involving anthropologists and sociologists, can provide a richer understanding of cultural dimensions in hateful content (Zhou et al., 2023). Such work can produce different forms of cultural features, such as switching pattern matrix (Bansal et al., 2020), bilingual language pairs (Jiang and Zubiaga, 2023), cross-cultural similarities (Arango et al., 2021), and social dynamics among users (Nozza, 2021).
@@ 7.4.4. Multilingual Pre-Trained Language Models
Multilingual PLMs can be applied to low-resource target languages without further training, thus bridging the data availability gap between high- and low-resource languages. However, there are significant computational costs during training and fine-tuning due to their large number of parameters (Pikuliak et al., 2021). Future research is geared towards optimising these multilingual PLMs for enhanced efficiency and interpretability. Additionally, two strategies are emerging to improve the generalisability and scalability of multilingual PLMs so that they can scale to handle large datasets or multiple languages simultaneously (Arango et al., 2022). The first emphasises pre-training models using data from relevant sources or language communities (such as social platforms and hateful domains), as demonstrated by XLM-T (Montariol et al., 2022) and AbuseXLMR (Gupta et al., 2022). This can mitigate model behaviour bias across different language communities. The second strategy focuses on models based on specific low-resource languages, dialects or those from similar language families, as seen with MuRIL for Indic languages (Das et al., 2022) and AraBERT for dialects (Husain and Uzuner, 2022), which requires intentional inclusion of dialectal and under-represented language data in both training and evaluation. Meanwhile, it is important to evaluate transfer effectiveness across both similar and distant language pairs, rather than relying solely on high-resource or typologically similar examples. Otherwise, multilingual PLMs risk reinforcing existing performance gaps, performing well for Indo-European languages while neglecting typologically distant or underrepresented languages.
@@ 7.4.5. Cross-lingual Training Strategies
Various cross-lingual training strategies have been proven effective. Adaptive training techniques can dynamically adjust to the specificities of different languages and dialects, as well as bridge the gap between offensive language detection and PLMs. By leveraging meta-learning or few-shot learning, models can be trained to quickly adapt to new low-resource languages with few labelled data (Mozafari et al., 2022). Furthermore, collaborative multi-task learning can make models more generalisable in cultural variations and resistant to overfitting by training them and sharing representations on multiple auxiliary tasks simultaneously across different languages or similar NLP tasks (Montariol et al., 2022). The use of adversarial training, where models are trained against adversarial examples, can also enhance their robustness and generalisation across languages. This is especially important for offensive language, as attackers often use subtle language tricks to obfuscate and bypass detection (Shi et al., 2022; Jin et al., 2020; Ryzhova et al., 2022). Moreover, cascade learning can greatly bolster model performance and enhance out-of-domain generalisability by initially training models on source data or more diverse datasets and subsequently fine-tuning the results by simultaneously or progressively incorporating target language data (Röttger et al., 2022a; Pelicon et al., 2021b). Besides, Vitiugin et al. (2021) highlight that human feedback can be of great value to help detect subtleties of abuse and phrases during model training, Ranasinghe and Zampieri (2021c) explore language-specific preprocessing like segmentation in morphologically rich languages such as Arabic and Turkish, and multi-source training, with augmented datasets, is also able to capture the diversities and commonalities of linguistic patterns across languages (Hande et al., 2021a; Markov et al., 2021; Chiril et al., 2019). Future research should focus on systematically comparing these strategies and identifying which are most effective under varying resource and language-family conditions.
@@ 7.4.6. Application of Large Language Models
With the emergence of Large Language Models (LLMs), the landscape of offensive language detection has rapidly shifted. Models such as GPT-3 (Brown et al., 2020), GPT-4 (OpenAI, 2023), Fine-tuned LAnguage Net (FLAN) (Wei et al., 2021) have demonstrated impressive multilingual capabilities, and recent surveys have highlighted their growing importance (Zhao et al., 2023). Instruction-tuned LLMs with prompting can already perform competitive detection across monolingual and multilingual settings without the need for task-specific fine-tuning (Plaza-del arco et al., 2023). Prompt design strategies, such as incorporating task descriptions or input-label demonstrations, further enhance performance (Chen et al., 2022; Han and Tang, 2022; Roy et al., 2023). LLMs also show promise in addressing data scarcity by generating synthetic examples for low-resource languages (Hartvigsen et al., 2022), and in providing explanations that improve interpretability (Wang et al., 2023). Despite their advantages, LLMs are not a full replacement for prior methods. They often underperform specialised multilingual PLMs on fine-grained tasks when little prompting expertise is available, and they remain costly to deploy at scale. Traditional approaches, such as leveraging domain-specific lexicons (e.g., HurtLex) or designing language-family-specific PLMs (e.g., MuRIL, AraBERT), still work well for targeted low-resource or domain-specific scenarios, offering more efficiency and control. A key direction is to explore how LLMs and prior CLTL techniques can be combined. For example, lexicon features and culturally-aware embeddings can be integrated as prompts or external knowledge sources to guide LLMs, while smaller multilingual PLMs can be lightweight alternatives in resource-constrained environments.
Overall, while early work on LLMs in this area was scarce at the time of our survey design, the rapid pace of development since then underscores their central role going forward. Future studies should explicitly compare LLM-based approaches with prior CLTL techniques to establish where LLMs truly offer an advantage and where traditional methods remain more effective.
@@ 8. Conclusion
Limitations
While this survey provides a comprehensive synthesis of CLTL-based offensive language detection, several limitations remain.
Focused scope: We focus exclusively on cross-lingual text classification for offensive language detection, excluding related modalities (e.g., multimodal detection) and other transfer scenarios such as cross-domain or cross-platform learning.
Coverage of literature: Our selection presents a comprehensive review of the literature which is updated up to the cut-off point of July 2023.
Benchmark comparison: Although we summarise performance trends, we do not conduct benchmarking experiments with reported results, which may vary due to evaluation setups.
Ethical Considerations
CLTL-based offensive language detection systems can raise several ethical considerations with the datasets, methods, and challenges we have discussed.
Bias in multilingual datasets: Multilingual offensive language datasets often suffer from topic bias, imbalanced language representation, and inconsistent annotation guidelines, which can propagate through CLTL models across languages and communities. Balanced and culturally diverse data collection, transparent reporting of dataset composition, and fairness-aware evaluation are needed.
Cultural sensitivity: Offensive expressions are often context-dependent, and models trained on data from one cultural context may misclassify content in another. Culturally-aware annotation and evaluation with native speakers from target communities are necessary.
Model transparency and interpretability: CLTL systems, particularly those based on PLMs, often act as“black boxes”. This lack of interpretability can hinder error analysis. Techniques such as attention or post-hoc visualisations (e.g., LIME, SHAP), and clear model training and evaluation setups can improve transparency.
Responsible use and risk mitigation: While CLTL-based systems can support harm reduction, they also carry risks of misuse for censorship, political suppression, or disproportionate moderation of certain language groups. Deployment should have clear governance frameworks, stakeholder engagement, and human-in-the-loop review mechanisms.
@@ Acknowledgements.
- Alakrot et al. (2018) Azalden Alakrot, Liam Murray, and Nikola S Nikolov. 2018. Dataset construction for the detection of anti-social behaviour in online communication in Arabic. Procedia Comp. Sci. 142 (2018), 174–181.
@@
- Aluru et al. (2021) Sai Saketh Aluru, Binny Mathew, Punyajoy Saha, and Animesh Mukherjee. 2021. A Deep Dive into Multilingual Hate Speech Classification. In Proc. of ECML PKDD. Springer, Cham, 423–439.
@@
- Bosco et al. ([n. d.]) Cristina Bosco, Dell’Orletta Felice, Fabio Poletto, Manuela Sanguinetti, Tesconi Maurizio, et al. [n. d.]. Overview of the evalita 2018 hate speech detection task. In Proc. of CEUR Workshop, Vol. 2263. CEUR, 1–9.
@@
- Chen et al. (2022) Yulong Chen, Yang Liu, Li Dong, Shuohang Wang, Chenguang Zhu, Michael Zeng, and Yue Zhang. 2022. AdaPrompt: Adaptive Model Training for Prompt-based NLP. In Findings of EMNLP. ACL, Abu Dhabi, UAE, 6057–6068.
@@
- Deng et al. (2022) Jiawen Deng, Jingyan Zhou, Hao Sun, Chujie Zheng, Fei Mi, Helen Meng, and Minlie Huang. 2022. COLD: A Benchmark for Chinese Offensive Language Detection. In Proc. of EMNLP. ACL, Abu Dhabi, UAE, 11580–11599.
@@
- Firmino et al. (2021) Anderson Almeida Firmino, Cláudio Souza de Baptista, and Anselmo Cardoso de Paiva. 2021. Using Cross Lingual Learning for Detecting Hate Speech in Portuguese. In Database and Expert Systems Applications. Springer, Cham, 170–175.
@@
- Jeong et al. (2022) Younghoon Jeong, Juhyun Oh, Jongwon Lee, Jaimeen Ahn, Jihyung Moon, Sungjoon Park, and Alice Oh. 2022. KOLD: Korean Offensive Language Dataset. In Proc. of EMNLP. ACL, Abu Dhabi, UAE, 10818–10833.
@@
- Lundberg and Lee (2017) Scott M Lundberg and Su-In Lee. 2017. A Unified Approach to Interpreting Model Predictions. In Proc. of NeurIPS 30. Curran Associates, Inc., 4765–4774.
@@
- Markov et al. (2021) Ilia Markov, Nikola Ljubešić, Darja Fišer, and Walter Daelemans. 2021. Exploring Stylometric and Emotion-Based Features for Multilingual Cross-Domain Hate Speech Detection. In Proc. of WASSA. ACL, Online, 149–159.
@@
- Mubarak et al. (2021) Hamdy Mubarak, Ammar Rashed, Kareem Darwish, Younes Samih, and Ahmed Abdelali. 2021. Arabic Offensive Language on Twitter: Analysis and Experiments. In Proc. of WANLP. ACL, Kyiv, Ukraine (Virtual), 126–135.
@@
- Ranasinghe and Zampieri (2021c) Tharindu Ranasinghe and Marcos Zampieri. 2021c. Multilingual Offensive Language Identification for Low-Resource Languages. ACM Transactions on Asian and Low-Resource Language Information Processing 21, 1 (2021), 13.
@@
- Rivaling (2021) Bilingual Word Embeddings Rivaling. 2021. Llod-driven bilingual word embeddings rivaling cross-lingual transformers in quality of life concept detection from french online health communities. In Proc. of SEMANTiCS, Vol. 53. 89.
@@
- Rodríguez et al. (2021) Sebastián E. Rodríguez, Héctor Allende-Cid, and Héctor Allende. 2021. Detecting Hate Speech in Cross-Lingual and Multi-lingual Settings Using Language Agnostic Representations. In Proc. of CIARP. Springer, Cham, 77–87.
- Rodr´ıguez-Sánchez et al. (2021) Francisco Rodríguez-Sánchez, Jorge Carrillo-de Albornoz, Laura Plaza, Julio Gonzalo, Paolo Rosso, Miriam Comet, and Trinidad Donoso. 2021. Overview of exist 2021: sexism identification in social networks. Procesamiento del Lenguaje Natural 67 (2021), 195–207.
- Rodr´ıguez-Sánchez et al. (2022) Francisco Rodríguez-Sánchez, Jorge Carrillo-de Albornoz, Laura Plaza, Adrián Mendieta-Aragón, Guillermo Marco-Remón, Maryna Makeienko, María Plaza, Julio Gonzalo, Damiano Spina, and Paolo Rosso. 2022. Overview of exist 2022: sexism identification in social networks. Procesamiento del Lenguaje Natural 69 (2022), 229–240.
@@
- Röttger et al. (2022a) Paul Röttger, Debora Nozza, Federico Bianchi, and Dirk Hovy. 2022a. Data-Efficient Strategies for Expanding Hate Speech Detection into Under-Resourced Languages. In Proc. of EMNLP. ACL, Abu Dhabi, UAE, 5674–5691.
- Röttger et al. (2022b) Paul Röttger, Haitham Seelawi, Debora Nozza, Zeerak Talat, and Bertie Vidgen. 2022b. Multilingual HateCheck: Functional Tests for Multilingual Hate Speech Detection Models. In Proc. of the WOAH Workshop. ACL, Seattle, Washington (Hybrid), 154–169.
@@
- Ryzhova et al. (2022) Anastasia Ryzhova, Dmitry Devyatkin, Sergey Volkov, and Vladimir Budzko. 2022. Training Multilingual and Adversarial Attack-Robust Models for Hate Detection on Social Media. Procedia Comp. Sci. 213 (2022), 196–202.
@@
- Stamou et al. (2022) Vivian Stamou, Iakovi Alexiou, Antigone Klimi, Eleftheria Molou, Alexandra Saivanidou, and Stella Markantonatou. 2022. Cleansing & expanding the HURTLEX(el) with a multidimensional categorization of offensive words. In Proc. of WOAH. ACL, 102–108.
@@
- Strus̈ et al. (2019) Julia Maria Strus̈, Melanie Siegel, Josef Ruppenhofer, Michael Wiegand, and Manfred Klenner. 2019. Overview of GermEval Task 2, 2019 shared task on the identification of offensive language. In Proc. of KONVENS. München, 352 – 363.
@@
- Zampieri et al. (2019) Marcos Zampieri, Shervin Malmasi, Preslav Nakov, Sara Rosenthal, Noura Farra, and Ritesh Kumar. 2019. SemEval-2019 Task 6: Identifying and Categorizing Offensive Language in Social Media (OffensEval). In Proc. of the SemEval Workshop. ACL, Minneapolis, Minnesota, USA, 75–86.